# Privy: Envisioning and Mitigating Privacy Risks for Consumer-facing AI Product Concepts

### v1

@@ 3.3.2. Facilitate privacy discussion and decision-making
While prior work shows privacy is often treated as an afterthought for compliance (Prybylo et al., 2024), participants appreciated how Privy beta enabled a more proactive review of AI product concepts. U3 reflected: “usually customers provide the requirement and we fulfill it, but now we can… initiate discussion with them based on potential privacy issues.” Others emphasized its alignment with the aspirational ‘Privacy by Design’ ideal, catching risks early and reducing downstream rework. As U6 explained: “That will save us a lot of time… we didn’t think about the privacy thing until the end… and then we had to spend another two weeks to work with other engineers.”
@@ 3.3.4. Emergent ideas: support mitigation and autonomy in privacy work
Participants also sought greater control over Privy beta-generated content to ensure it aligned with their work. Some modified or expanded Privy beta’s outputs to better capture product nuances, while others critiqued the outputs for falling short of their professional standards: “it didn’t think outside the box as much as like a professional would” (U10).
@@ 4. Privy
We distilled five design goals (DG1–DG5) to guide the creation of Privy, a tool that supports AI practitioners in envisioning and mitigating privacy risks for novel AI product concepts. Privy leads users through a structured workflow: they first enter a natural language description of the product and its data sources; then outline use cases and associated beneficiaries; specify the AI capabilities and requirements to support those use-cases; and, finally, identify the most relevant and severe risks from the AI privacy taxonomy (Lee et al., 2024b) based on these capabilities and requirements, use-cases, and the product description. We implemented two versions of Privy, modeled after prior AI risk assessment approaches (Section 2.3): (1) Privy-LLM, an interactive system that leverages LLMs for human–AI collaboration, and (2) Privy-Template, a structured worksheet. Both embody the same scaffolded workflow. In the following sections, we use Privy-LLM’s interfaces to illustrate how we achieve each design goal, and highlight LLM-powered features that support human–AI collaboration. We present the Privy-Template’s interface in Section 4.5.
@@ 4.1. DG1: Help practitioners elicit the AI capabilities and requirements that entail privacy risks
Our formative study showed that practitioners’ ability to identify privacy risks is grounded in understanding what their AI product does and what it requires to work. Likewise, prior work finds that practitioners feel more accountable for harms when reasoning about how their own products may cause them, rather than abstracting harms (Widder and Nafus, 2023). To that end, Privy helps practitioners articulate capabilities and requirements of their AI product concepts grounded in their envisioned use cases (Section 4.1.1); Privy-LLM augments this process with LLM-powered suggestions (Section 4.1.2).
@@ 4.1.1. AI Capability & Requirement Scaffolder
Privy guides users through three steps: (1) describing their AI product concepts, (2) brainstorming use cases, and (3) summarizing resulting AI capabilities and requirements.
@@ 4.1.2. Human-AI Collaboration in Scaffolding AI Capabilities and Requirements
For Privy-LLM, we leverage an LLM to help users brainstorm more use cases and summarize their AI products’ capabilities and requirements. For this paper, we specifically used OpenAI’s GPT-4.1 model, though Privy could work with any back-end LLM that supports the Chat Completions API.
@@
From the product description, the LLM suggests intended and unintended use cases with the corresponding beneficiaries or exploiters. Each use case is shown as a sentence describing how a user could use the AI product in a specific context — e.g., “a user can use a textbook tutor AI to generate a list of recommended readings based on prior coursework”. To diversify generated use cases, we instruct the LLM to consider both high-stakes (e.g., college entrance exams) and low-stakes (e.g., take-home quizzes) contexts. Privy-LLM initially offers four examples (intended/unintended × high-/low-stakes) and allows users to request more (Fig. 1B-2).
@@
We prompt LLMs to extract an AI product’s capabilities using the AI capabilities taxonomy (Yildirim et al., 2023), and to infer associated requirements based on data collection, processing, dissemination, and infrastructure needs (Lee et al., 2024b). First, the LLM generates capability–requirement pairs from each envisioned use case. For example, from the use case “a user requests recommended readings based on prior coursework”, Privy-LLM derives the capability to “generate personalized reading recommendations and factual responses by analyzing textbook content and students’ historical coursework”, along with the following requirements: “(1) collect textbook materials and student learning histories, (2) process and analyze these data to generate recommendations, (3) disseminate recommendations to users, and (4) integrate with learning platforms for user interaction and data exchange”. A second LLM then merges all pairs into a concise summary, shown in Privy-LLM’s summary node (Fig. 1C).
@@ 4.2.2. Human-AI Collaboration in Envisioning Privacy Risks
Privy-LLM generates examples of privacy risks for users’ AI product concepts, with the associated stakeholders, tailored to the product’s description, use cases, and AI capabilities and requirements. For example, for a textbook tutor AI, Privy-LLM may suggest a surveillance risk such as “monitoring individuals’ learning histories and performance”, particularly for “students who use the platform”. Privy-LLM produces one example per taxonomy risk type (12 in total) (Lee et al., 2024b), which users can review and selectively incorporate into their assessments (Fig. 1D-2).
@@ 4.3.2. Human-AI Collaboration in Brainstorming Mitigation Strategies
To support ideation, Privy-LLM generates open-ended questions tailored to users’ AI product concepts, encouraging users to consider how their systems might incorporate privacy-preserving features or practices in response to specific risks. For instance, if a textbook tutor AI raises surveillance risks due to monitoring students’ learning history, a mitigation prompt might encourage reflection on “how to provide users with tools to control, limit, or delete collected data”.
@@ 4.4.1. Privacy Risk Summarizer
Privy collates outputs from the capability & requirement scaffolder, risk explorer, and risk mitigator into a shareable privacy impact assessment report. The interface is modeled after an established AI impact assessment report template (Bogucka et al., 2024b) and presents three sections (Appendix A): (1) product description, use cases, and AI capabilities & requirements; (2) identified privacy risks; and (3) proposed mitigation strategies.
@@ 4.5.1. User Agency in Human-AI Collaboration
Figure 4. Privacy risk assessment panel: users can edit the node content to capture their own perspectives, including the risk description, impacted stakeholders, and their assessment of the identified risk’s relevancy and severity
@@
To promote critical integration of LLM outputs (Lee et al., 2025), we embed design friction into Privy-LLM. For example, users must manually assess the relevance and severity of AI-suggested risks before including them (Section 4.2.1). Similarly, when brainstorming mitigations, users receive provocative questions rather than concrete suggestions — encouraging them to generate context-appropriate solutions themselves (Section 4.3.2).
4.5.2. Privy-Template: Privy as a structured worksheet
To isolate the role of LLM assistance, we developed Privy-Template, a static version of Privy that follows the same workflow (DG1–DG4) without LLM assistance. Delivered as a structured Google Doc, it includes the full AI privacy taxonomy (Lee et al., 2024b) (Fig. 5A) and guides users through three sections: Product Information, Privacy Risks, and Mitigation Strategies (Fig. 5; full version in Appendix B).
Figure 5. Privy-Template: the worksheet is administered in the form of a structured Google Doc with three main sections: Product Information, Privacy Risks, and Mitigation Strategies. The worksheet contains (A) the AI privacy taxonomy (Lee et al., 2024b) in a detailed table as a reference. See the full worksheet in Appendix B.
For mitigation, Privy-Template substitutes dynamic LLM-generated provocations with static ones grounded in the SPAF (Das et al., 2022): How can the product be designed to 1) enhance users’ awareness of the risk? 2) enhance users’ motivation to address the risk? 3) provide users with the ability to manage the risk? Privy-Template builds on an established AI impact assessment report template (Bogucka et al., 2024b), ensuring usability and shareability (DG4).
In our evaluation, comparing Privy-LLM and Privy-Template allowed us to assess how LLM assistance affected the quality of privacy assessments and practitioners’ perceived usefulness of Privy.
4.6. Privy-LLM Implementation
We implement Privy-LLM as a web client: we implement the interactive front-end with React.js, and use a Python Flask server as the backend. We leverage React Flow [^1] https://github.com/xyflow/xyflow/tree/v11 and Material UI [^2] https://github.com/mui/material-ui to implement node-link interactions and styling. Finally, we leverage a large language model, GPT-4.1, to enable the human-AI collaboration we described, and follow best practices when designing our prompts, including the use of few-shot learning and chain-of-thought prompting (Wei et al., 2022).
@@ 5. Evaluation User Study
To address our three research questions (Section 1), we conducted a between-subjects study with $N=24$ industry AI practitioners who were not privacy experts but worked across diverse AI product roles. Participants were randomly assigned to use either Privy-LLM ($N_{llm}=12$) or Privy-Template ($N_{template}=12$) to create privacy impact assessments for two pre-defined AI product concepts. We also recruited $M=13$ privacy experts to independently evaluate the resulting reports; these experts were not told which reports were produced by Privy-LLM or Privy-Template. Qualitatively (RQ1–RQ3), we analyzed participants’ think-aloud and interview transcripts. Quantitatively (RQ2), we compared practitioners’ self-report perceptions and expert-rated report quality across conditions (Privy-LLM vs. Privy-Template).
@@ 5.1. Participants
We conducted four pilot studies with Privy-LLM and four with Privy-Template (eight total). Each study session lasted 90 minutes, and participants received a \$150 Amazon gift card as compensation. Of the 24 participants, 17 identified as men and seven as women, spanning a range of industry roles (Appendix D.1). None were privacy experts: four reported being knowledgeable about privacy, 15 had some knowledge, four had limited knowledge, and one had none.
@@ 5.2. Study Design
We conducted one-on-one remote sessions over Zoom, recording video, computer screens, and audio for analysis. Participants were randomly assigned to one of four conditions: using either Privy-LLM or Privy-Template to assess one of two pre-defined AI product concepts (Table 1), with six participants per condition.
@@
Privy-LLM & AI Meeting Assistant P1, P2, P3, P4, P5, P6
Privy-LLM & Dynamic Audience Selection P7, P8, P9, P10, P11, P12
Privy-Template & AI Meeting Assistant P13, P14, P15, P16, P17, P18
Privy-Template & Dynamic Audience Selection P19, P20, P21, P22, P23, P24
@@ 5.2.1. Privacy Impact Assessment (PIA) Activity
During risk identification, participants described three risks — from the twelve total in the AI privacy taxonomy (Lee et al., 2024b) — that they felt the most important to mitigate. Next, during risk mitigation, they proposed alternative designs of the AI product concept or other strategies (e.g., legal/policy reviews, data access controls) that they would like to share with their team to mitigate the three identified risks. Participants were instructed to work entirely within Privy’s features and interfaces, and to think-aloud throughout. Participants were also asked to reflect on their approach, once after completing risk identification and once after completing risk mitigation. Finally, we reviewed the final report — including the identified risks and proposed mitigation plan — together with the participants for any final adjustments before proceeding to the post-activity interview.
@@ 5.2.2. Post-activity Semi-structured Interviews
Following the PIA activity, participants re-rated the AI product concept’s benefit and intrusiveness, and completed survey items (Appendix D.2.2) on the tool’s usefulness (Privy design goal alignment) and usability (System Usability Scale (Brooke et al., 1996)). Semi-structured interviews (Appendix D.2.1) probed how Privy influenced their approach to privacy, tool feedback, and broader challenges in conducting PIAs.
@@ 5.2.3. Privacy Impacts Assessment Report Evaluation by Privacy Experts
We collected a total of 24 PIA reports comprising 72 identified privacy risks (with product-specific descriptions of those risks) and 24 mitigation plans. To evaluate the quality of these reports, we recruited 13 external privacy experts (ten industry, three academia) through our networks. All self-identified as privacy experts, with roles spanning researchers, engineers, consultants, and a professor. None had prior exposure to Privy.
Each expert reviewed four PIA reports for the same product concept (AI Meeting Assistant or Dynamic Audience Selection), along with the accompanying product description (Table 2). We originally aimed for 12 experts, each evaluating four reports (two from Privy-LLM, two from Privy-Template), such that every report was evaluated twice. Through our parallelized recruitment outreach, we ultimately recruited an additional privacy expert, resulting in four reports being evaluated by three experts [^3] For the additional privacy expert evaluation (E13), we used the same procedure to randomly sample four assessment reports (two from Privy-LLM, two from Privy-Template) of the AI Meeting Assistant product concept.. Experts were blinded to the study conditions (Privy-LLM vs. Privy-Template), and the reports they reviewed were randomly and evenly sampled.

### v2

@@ 3.3.2. Facilitate privacy discussion and decision-making
While prior work shows privacy is often treated as an afterthought for compliance (Prybylo et al., 2024), participants appreciated how Privybeta enabled a more proactive review of AI product concepts, and extrapolated its impacts on their product development. U3 reflected: “usually customers provide the requirement and we fulfill it, but now we can… initiate discussion with them based on potential privacy issues.” Others emphasized its alignment with the aspirational ‘Privacy by Design’ ideal, catching risks early and reducing downstream rework. As U6 explained: “That will save us a lot of time… we didn’t think about the privacy thing until the end… and then we had to spend another two weeks to work with other engineers.”
@@ 3.3.4. Emergent ideas: support mitigation and autonomy in privacy work
Participants also sought greater control over Privybeta-generated content to ensure it aligned with their work. Some modified or expanded Privybeta’s outputs to better capture product nuances, while others critiqued the outputs for falling short of their standards for work (e.g., novelty, creativity): “it didn’t think outside the box as much as like a professional would” (U10).
@@ 4. Privy
From our formative study findings (Section 3.3), we distilled five design goals (DG1–DG5) to guide the creation of Privy, a tool that supports AI practitioners in envisioning and mitigating privacy risks for novel AI product concepts. Privy leads users through a structured workflow: they first enter a natural language description of the product and its data sources; then outline use cases and associated beneficiaries; specify the AI capabilities and requirements to support those use-cases; and, finally, identify the most relevant and severe risks from the AI privacy taxonomy (Lee et al., 2024b) based on these capabilities and requirements, use-cases, and the product description. We implemented two versions of Privy, modeled after prior AI risk assessment approaches (Section 2.3): (1) PrivyLLM, an interactive system that leverages LLMs for human–AI collaboration, and (2) PrivyTemplate, a structured worksheet. Both embody the same scaffolded workflow. In the following sections, we use PrivyLLM’s interfaces to illustrate how we achieve each design goal, and highlight LLM-powered features that support human–AI collaboration. We present the PrivyTemplate’s interface in Section 4.5.
@@ 4.1. DG1: Help practitioners elicit the AI capabilities and requirements that entail privacy risks
Our formative study showed that practitioners’ ability to identify privacy risks is grounded in understanding what their AI product does and what it requires to work. Likewise, prior work finds that practitioners feel more accountable for harms when reasoning about how their own products may cause them, rather than abstracting harms (Widder and Nafus, 2023). To that end, Privy helps practitioners articulate capabilities and requirements of their AI product concepts grounded in their envisioned use cases (Section 4.1.1); PrivyLLM augments this process with LLM-powered suggestions (Section 4.1.2).
@@ 4.1.1. AI Capability & Requirement Scaffolder
Both PrivyTemplate and PrivyLLM guide users through three steps: (1) describing their AI product concepts, (2) brainstorming use cases, and (3) summarizing resulting AI capabilities and requirements.
@@ 4.1.2. Human-AI Collaboration in Scaffolding AI Capabilities and Requirements
For PrivyLLM, we leverage an LLM to help users brainstorm more use cases and summarize their AI products’ capabilities and requirements (see Supplementary Material for LLM prompts). For this paper, we specifically used OpenAI’s GPT-4.1 model, though Privy could work with any back-end LLM that supports the Chat Completions API.
@@
From the product description, the LLM suggests intended and unintended use cases with the corresponding beneficiaries or exploiters. Each use case is shown as a sentence describing how a user could use the AI product in a specific context — e.g., “a user can use a textbook tutor AI to generate a list of recommended readings based on prior coursework”. To diversify generated use cases, we instruct the LLM to consider both high-stakes (e.g., college entrance exams) and low-stakes (e.g., take-home quizzes) contexts. PrivyLLM initially offers four examples (intended/unintended × high-/low-stakes) and allows users to request more (Fig. 1B-2).
@@
We prompt LLMs to extract an AI product’s capabilities using the AI capabilities taxonomy (Yildirim et al., 2023), and to infer associated requirements based on data collection, processing, dissemination, and infrastructure needs (Lee et al., 2024b). First, the LLM generates capability–requirement pairs from each envisioned use case. For example, from the use case “a user requests recommended readings based on prior coursework”, PrivyLLM derives the capability to “generate personalized reading recommendations and factual responses by analyzing textbook content and students’ historical coursework”, along with the following requirements: “(1) collect textbook materials and student learning histories, (2) process and analyze these data to generate recommendations, (3) disseminate recommendations to users, and (4) integrate with learning platforms for user interaction and data exchange”. A second LLM then merges all pairs into a concise summary, shown in PrivyLLM’s summary node (Fig. 1C).
@@ 4.2.1. Privacy Risk Explorer
Screenshot of the risk ranking panel, which displays privacy risks identified by the user in orange boxes on the left. Each box includes a risk description, risk type, and relevancy and severity tags. Using a drag-and-drop interaction, users move risks into a ranked list on the right.
@@ 4.2.2. Human-AI Collaboration in Envisioning Privacy Risks
We use a few-shot prompting approach (Brown et al., 2020) to generate examples of privacy risks for users’ AI product concepts, with the associated stakeholders, tailored to the product’s description, use cases, and AI capabilities and requirements. For example, for a textbook tutor AI, PrivyLLM may suggest a surveillance risk such as “monitoring individuals’ learning histories and performance”, particularly for “students who use the platform”. PrivyLLM produces one example per privacy risk type defined in Lee et al. (Lee et al., 2024b) (12 in total), which users can review and selectively incorporate into their assessments (Fig. 1D-2).
@@ 4.3.1. Privacy Risk Mitigator
Screenshot of the risk mitigation interface, which guides users through coming up with strategies to address identified privacy risks. On the left, a selected surveillance risk card is displayed in an orange box with its risk description, risk type, and relevancy and severity score tags. Below, users are prompted to brainstorm mitigation strategies by adding ideas to a “Full Mitigation Plan” on the right. Users can click a button for suggested brainstorming questions generated by Privy-LLM, phrased as guiding questions. A confidence rating scale from Strongly agree to Strongly disagree (5-point Likert scale) follows, where users indicate how strongly they agree with the statement “I’m confident with the mitigation strategies in addressing this risk.” On the right, the “Full Mitigation Plan” is shown in a blue box, listing users’ drafted strategies.
@@ 4.3.2. Human-AI Collaboration in Brainstorming Mitigation Strategies
To support ideation, PrivyLLM generates open-ended questions tailored to users’ AI product concepts, encouraging users to consider how their systems might incorporate privacy-preserving features or practices in response to specific risks. For instance, if a textbook tutor AI raises surveillance risks due to monitoring students’ learning history, a mitigation prompt might encourage reflection on “how to provide users with tools to control, limit, or delete collected data”.
@@ 4.4.1. Privacy Risk Summarizer
Privy collates outputs from the capability & requirement scaffolder, risk explorer, and risk mitigator into a shareable privacy impact assessment report. The interface is modeled after an established AI impact assessment report template (Bogucka et al., 2024b) and presents three sections (Appendix A): (1) product description, use cases, and AI capabilities & requirements; (2) identified privacy risks along with relevancy and severity ratings provided by practitioners; and (3) proposed mitigation strategies.
@@ 4.5.1. User Agency in Human-AI Collaboration
Figure 4. Privacy risk assessment panel: users can edit the node content to capture their own perspectives, including the risk description, impacted stakeholders, and their assessment of the identified risk’s relevancy and severity.
Interface for editing a privacy risk node. A card describes a Surveillance Risk for an AI meeting assistant, labeled with medium relevancy and severity, and includes a button to “View Impacted Stakeholders.” Clicking this button expands a panel where users can adjust relevancy and severity ratings and add or edit impacted stakeholders. On the left, selecting the edit icon opens an inline panel that allows direct modification of the risk description.
@@
To promote critical integration of LLM outputs (Lee et al., 2025), we embed design friction into PrivyLLM. For example, users must manually assess the relevance and severity of AI-suggested risks before including them (Section 4.2.1). Similarly, when brainstorming mitigations, users receive provocative questions rather than concrete suggestions — encouraging them to generate context-appropriate solutions themselves (Section 4.3.2).
4.5.2. PrivyTemplate: Privy as a structured worksheet
To isolate the role of LLM assistance, we developed PrivyTemplate, a static version of Privy that follows the same workflow (DG1–DG4) without LLM assistance. Delivered as a structured Google Doc, it includes the full AI privacy taxonomy (Lee et al., 2024b) (Fig. 5A) and guides users through three sections: Product Information, Privacy Risks, and Mitigation Strategies (Fig. 5; full version in Appendix B).
Figure 5. PrivyTemplate: the worksheet is administered in the form of a structured Google Doc with three main sections: Product Information, Privacy Risks, and Mitigation Strategies. The worksheet contains (A) the AI privacy taxonomy (Lee et al., 2024b) in a detailed table as a reference. See the full worksheet in Appendix B.
Interface of Privy-Template for documenting product information, identifying privacy risks, and mitigation strategies. Product information fields include use cases, AI capabilities and requirements. Risks are identified and assessed. The mitigation section captures strategies and confidence ratings. A reference panel lists types and definitions of AI privacy risks.
For mitigation, PrivyTemplate substitutes dynamic LLM-generated provocations with static ones grounded in the SPAF (Das et al., 2022): How can the product be designed to 1) enhance users’ awareness of the risk? 2) enhance users’ motivation to address the risk? 3) provide users with the ability to manage the risk? PrivyTemplate builds on an established AI impact assessment report template (Bogucka et al., 2024b), ensuring usability and shareability (DG4).
In our evaluation, comparing PrivyLLM and PrivyTemplate allowed us to assess how LLM assistance affected the quality of privacy assessments and practitioners’ perceived usefulness of Privy.
4.6. PrivyLLM Implementation
We implement PrivyLLM as a web client: we implement the interactive front-end with React.js, and use a Python Flask server as the backend. We leverage React Flow [^2] https://github.com/xyflow/xyflow/tree/v11 and Material UI [^3] https://github.com/mui/material-ui to implement node-link interactions and styling. We leverage a large language model, GPT-4.1, to enable the human-AI collaboration features we described, and follow best practices when designing our prompts, including the use of few-shot learning and chain-of-thought prompting (Wei et al., 2022). For the few-shot examples, we manually curated 23 different AI product concepts across ten types of AI technology  [^4] The ten types of AI technologies are: computer vision, generative AI, speech-voice system, decision support system, information retrieval, natural language processing, user modeling, predictive analytics, recommendation system, and conversational AI., referenced on real-world AI products sourced from the AIAAIC database (Pownall, 2023). We included our LLM prompts, few-shot examples, and additional LLM-generated examples (e.g., use cases, privacy risks) in the Supplementary Materials.
@@ 5. Evaluation User Study
To address our three research questions (Section 1), we conducted a between-subjects study with $N=24$ industry AI practitioners who were not privacy experts but worked across diverse AI product roles. Participants were randomly assigned to use either PrivyLLM ($N_{llm}=12$) or PrivyTemplate ($N_{template}=12$) to create privacy impact assessments for two pre-defined AI product concepts. We also recruited $M=13$ privacy experts — who work as privacy practitioners and/or researchers in industry or academia — to independently evaluate the resulting reports; these experts were not told which reports were produced by PrivyLLM or PrivyTemplate. Qualitatively (RQ1–RQ3), we analyzed participants’ think-aloud and interview transcripts. Quantitatively (RQ2), we compared practitioners’ self-report perceptions and expert-rated report quality across conditions (PrivyLLM vs. PrivyTemplate).
@@ 5.1. Participants
We conducted four pilot studies with PrivyLLM and four with PrivyTemplate (eight total). Each study session lasted 90 minutes, and participants received a \$150 Amazon gift card as compensation. Of the 24 participants, 17 identified as men and seven as women, spanning a range of industry roles (Appendix D.1). None were privacy experts: four reported being knowledgeable about privacy, 15 had some knowledge, four had limited knowledge, and one had none.
@@ 5.2. Study Design
We conducted one-on-one remote sessions over Zoom, recording video, computer screens, and audio for analysis. Participants were randomly assigned to one of four conditions: using either PrivyLLM or PrivyTemplate to assess one of two pre-defined AI product concepts (Table 1), with six participants per condition.
@@
PrivyLLM & AI Meeting Assistant P1, P2, P3, P4, P5, P6
PrivyLLM & Dynamic Audience Selection P7, P8, P9, P10, P11, P12
PrivyTemplate & AI Meeting Assistant P13, P14, P15, P16, P17, P18
PrivyTemplate & Dynamic Audience Selection P19, P20, P21, P22, P23, P24
@@ 5.2.1. Privacy Impact Assessment (PIA) Activity
During risk identification, participants described three risks — from the twelve total in the AI privacy taxonomy (Lee et al., 2024b) — that they felt the most important to mitigate. Next, during risk mitigation, they proposed alternative designs of the AI product concept or other strategies (e.g., legal/policy reviews, data access controls) that they would like to share with their team to mitigate the three identified risks. Participants were instructed to work entirely within Privy’s features and interfaces, and to think-aloud throughout. Participants were also asked to reflect on their approach, once after completing risk identification and once after completing risk mitigation. Finally, we reviewed the final report — including the identified risks and proposed mitigation plan — together with the participants for any final adjustments before proceeding to the post-activity interview. We include the privacy risks identified and the proposed mitigation plans from each participant in the Supplementary Materials.
@@ 5.2.2. Post-activity Semi-structured Interviews
Following the PIA activity, participants re-rated the AI product concept’s benefit and intrusiveness, and completed survey items (Appendix D.2.2) on the tool’s usefulness (Privy design goal alignment) and usability (System Usability Scale (Brooke and others, 1996)). Semi-structured interviews (Appendix D.2.1) probed how Privy influenced their approach to privacy, tool feedback, and broader challenges in conducting PIAs.
@@ 5.2.3. Privacy Impacts Assessment Report Evaluation by Privacy Experts
We collected a total of 24 PIA reports comprising 72 identified privacy risks (with product-specific descriptions of those risks) and 24 mitigation plans. To evaluate the quality of these reports, we recruited 13 external privacy experts (ten industry, three academia) through our networks. All self-identified as privacy experts, with roles spanning researchers, privacy engineers, privacy program managers, consultants, and a professor. None had prior exposure to Privy.
Each expert reviewed four PIA reports for the same product concept (AI Meeting Assistant or Dynamic Audience Selection), along with the accompanying product description (Table 2). We originally aimed for 12 experts, each evaluating four reports (two from PrivyLLM, two from PrivyTemplate), such that every report was evaluated twice. Through our parallelized recruitment outreach, we ultimately recruited an additional privacy expert, resulting in four reports being evaluated by three experts [^5] For the additional privacy expert evaluation (E13), we used the same procedure to randomly sample four assessment reports (two from PrivyLLM, two from PrivyTemplate) of the AI Meeting Assistant product concept.. Experts were blinded to the study conditions (PrivyLLM vs. PrivyTemplate), and the reports they reviewed were randomly and evenly sampled.